An electric coil marked 120V is used to supply current to melt a solid at its melting point for 10 minutes. If the original temperature of the solid is -50C below its melting point and its mass is 0.5Kg, what is the specific latent heat of fusion? (Cw=4200J/Kg/K, Cs= 2.2x10⁴J/Kg/K)
1
Expert's answer
2021-09-08T09:45:02-0400

Gives

E=120V

Time=10 minute=600 sec

Temperature (T)=-50°c

Mass(m)=0.5kg

Assume i=1Amp

We know that

specific latent heat of fusion

"L=\frac{E}{m}\rightarrow(1)"

"P=\frac{E}{t}\\E=P\times t\rightarrow(2)"

Equation (1) and equation (2) we can be written as


"L=\frac{tP}{m}=\frac{t\times i V}{m}"

Put value

"L=\frac{t\times iV}{m}\\L=\frac{120\times 1\times 600}{0.5}=144000 J/K"